OpenStudy (anonymous):

If f(x) is strictly concave up, then f '(a) < (f(b) – f(a))/(b – a) for a < b. true or false?

OpenStudy (anonymous):

TRUE lets check \[f(x)=x^3\]is strictly increasing  if we take a=0,and a<b let b=1 \[f^'(0)=3(0)^2=0\] the average gradient between a and b \[=\frac{ f(b)-f(a) }{ b-a }\] is\[\frac{ 1^3-0^3 }{ 1-0 }=1\] \[f^'(a)<average, gradient\],if we try again for x=1 and b=2 \[f^'(1)=3\]\[avg=\frac{ 2^3-1^3 }{ 2-1 }=7\] again 3<7 true

OpenStudy (noelgreco):

Yours is strictly increasing, not concave up.

OpenStudy (stacey):

 Hopefully, you can see by the graph and the two examples that the slope of f(a) will always be less than the slope of a line from a point at x=a to a point at x=b (where a<b)